How long does it take a packet of 500 bytes length to propagate over a link of distance 1,500 km, propagation speed 2.5 x 10^8 m/s, and transmission rate 5 Mbps? More generally, how long does it take a packet of length L to propagate over a link of distance d, propagation speed s, and transmission rate R bps? Does this delay depend on packet length? Does this delay depend on transmission rate? Explain your choices.
Answer #1

Delays in Packet switching :

  1. Transmission Delay
  2. Propagation Delay
  3. Queuing Delay

Transmission Delay: the time required to put data bits on the wire/communication medium. It depends on the length of the packet and bandwidth of the network.

 

Transmission delay = L/R

= 8 bits/byte * 500 bytes / 5,000,000 bps

= 0.8 ms

Propagation Delay: the time required for bits to reach the destination from the start point. Factors on which the Propagation delay depends are Distance and propagation speed.

Propagation delay =distance/transmission speed = d/s

= 1,500 x10^3/ 2.5×10^8

= 6 ms

Queuing Delay :
Queuing delay is the time a job waits in a queue until it can be executed. It depends on congestion. It is the time difference between when the packet arrived Destination and when the packet data was processed or executed.

Average Queuing delay = (N-1)L/(2*R)=0(queue is iniatly empty).

Therefore, the total time =0. 8ms + 6ms = 6.8 ms

The propagation delay will depend only on distance and speed.

But, transmission delay depends on packet length and bandwidth. Hence, the total delay depends on the packet length and transmission rate.